A. You have just won the Strayer Lottery jackpot of $11,000,000. You will be paid in 26 equal annual installments beginning immediately. If you had the money now, you could invest it in an account with a quoted annual interest rate of 9% with monthly compounding of interest. What is the present value of the payments you will receive?

B. In your own words and using various bond websites, please locate one of each of the following bond ratings: AAA, BBB, CCC, and D. Please describe the differences between the bond ratings. Identify the strengths and weaknesses of each rating.
Answer #1

A. The winning amount = $1100000

Amount of each installment = $1100000/26 = $423077

Interest rate is 9% annually which is compounded monthly,

r = 9/12 = .0075, for monthly compounding interest

n = 26*12 = 312 months

Therefore, present value that you will receive = 423077 + 423077 {[1-(1+0.0075)^-(312-1)]/0.0075} = $51310712